OpenStudy (anonymous):

what is the approximate perimeter of a semicircle with a base of 5? PLEASE HELP ASAP

OpenStudy (anonymous):

Circumference of a circle is \( \pi d \) , where d=5 a semicircle has a perimeter that is half of the circle, so divide it by two \(\Large\frac{\pi d}{2}\)

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

no problem :)